OpenStudy (anonymous):

What are quartiles! Example problem: What is the first quartile (Q1) of the data set? 51, 42, 46, 53, 66, 70, 90, 70

OpenStudy (amistre64):

quartiles are positions of the data set that are divsions of 4

OpenStudy (anonymous):

so do you....add up the data then divide by 4? Seems logical right??

OpenStudy (anonymous):

idk....

OpenStudy (amistre64):

51, 42, 46, 53, 66, 70, 90, 70 | | | q1 q2 q3 notice that when we divide the set into 4 parts, the divisions end up between 2 values so the q values are the averages

OpenStudy (anonymous):

so q1 is....46.5?

OpenStudy (amistre64):

well, not adding up the values, but rather, dividing the number of data points by 4 there are 8 numbers, 8/4 = 2, this is a basic quartile interval for this set. Q1 is: 42+46 = 88/2 = 44

OpenStudy (amistre64):

spose we have an odd number of data points, so lets modify this by adding one new data point, say 73 51, 42, 46, 53, 66, 70, 90, 70, 73 we would have 9 data points; 9/4 would be the basic quartile setup Q1 = 9/4 times 1 = 2.25 ... since we have a decimal, we round up to 3. Q1 is the 3rd data point, 46 Q2 = 9/4 times 2 = 4.5 ... since we have a decimal, we round up to 5. Q2 is the 5th data point, 66

OpenStudy (amistre64):

since we have an even number of data points, we get divisions that have no decimals; and therefore we take the average of 2 data points ... 51, 42, 46, 53, 66, 70, 90, 70 we have 8 data points; 8/4 = 2 would be the basic quartile setup Q1 = 2 times 1 = 2 ... since we dont have a decimal, we round up to average the 2nd and 3rd data points, Q1 = (42+46)/2 Q2 = 2 times 2 = 4 ... since we dont have a decimal, we round up to average the 4th and 5th data points, Q2 = (66+70)/2 etc .....

OpenStudy (anonymous):

Ahhhh......

OpenStudy (amistre64):

i forgot to edit out the "round up" parts :) if we dont have a decimal, we average the 2 data points

OpenStudy (anonymous):

Oh....ok thx!

OpenStudy (amistre64):

good luck :) percentiles are the same concept, except for a divide by 100 instead of a divide by 4